# Slaying Axion-Like Particles via Gravitational Waves and Primordial Black Holes from Supercooled Phase Transition

### v1

@@ Abstract
We study the formation of primordial black holes (PBHs) from density fluctuations due to supercooled phase transitions (PTs) triggered in an axion-like particle (ALP) model. We find that the mass of the PBHs is inversely correlated with the ALP decay constant $f_{a}$. For instance, for $f_{a}$ varying from ${\cal O}$(100 MeV) to ${\cal O}$($10^{12}$ GeV), the PBH mass varies between $(10^{3}-10^{-24})M_{\odot}$. We then identify the ALP parameter space where the PBH can account for the entire (or partial) dark matter fraction of the Universe, in a single (multi-component) dark matter scenario, with the ALP being the other dark matter candidate. The PBH parameter space ruled out by current cosmological and microlensing observations can thus be directly mapped onto the ALP parameter space, thus providing new bounds on ALPs, complementary to the laboratory and astrophysical ALP constraints. Similarly, depending on the ALP couplings to other Standard Model particles, the ALP constraints on $f_{a}$ can be translated into a lower bound on the PBH mass scale. Moreover, the supercooled PT leads to a potentially observable stochastic gravitational wave (GW) signal at future GW observatories, such as aLIGO, LISA and ET, that acts as another complementary probe of the ALPs, as well as of the PBH dark matter. Finally, we show that the recent NANOGrav signal of stochastic GW in the nHz frequency range can be explained in our model with $f_{a}\simeq(10~{}{\rm GeV}-1~{}{\rm TeV})$.
@@ 1 Introduction
The first direct detection of gravitational waves (GWs) by the LIGO-VIRGO collaboration [1] has opened up new avenues to explore the Universe. The known astrophysical sources of GWs can be broadly split into three categories [2]: (i) transient signals (with a duration between a millisecond and several hours) emitted by the merger of two compact objects, like black hole or neutron star binaries, or supernova core collapse; (ii) long-duration (or steady-state) signals, e.g. from spinning neutron stars or from binary white-dwarf mergers; and (iii) stochastic background arising from the superposition of unresolved astrophysical sources. Stochastic GW background (SGWB) is also a unique probe of the early Universe, as the Universe is transparent to GWs right from the wee moments of the Big Bang, unlike other cosmic relics like photons and neutrinos. Although LIGO-VIRGO has only set an upper limit on the SGWB [3, 4, 5], the increased sensitivity of future GW detectors in a wide frequency range from nHz-kHz, such as SKA [6], GAIA/THEIA [7], MAGIS [8], AION [9], AEDGE [10], $\mu$ARES [11], LISA [12], TianQin [13], Taiji [14], DECIGO [15], BBO [16], ET [17], CE [18], as well as recent proposals for high-frequency GW searches in the MHz-GHz regime [19, 20, 21, 22], makes the future detection of SGWB an exciting real possibility. The recent evidence supporting the existence of a SGWB coming from several Pulsar Timing Arrays (PTAs), namely, NANOGrav [23], EPTA [24], PPTA [25] and CPTA [26], has added fuel to the excitement and opened a floodgate of papers with various interpretations, from mundane astrophysics to exotic new physics; see e.g. Refs. [27, 28, 29].
@@
Interestingly, the FOPT energy scale LIGO is sensitive to roughly coincides with the lowest possible energy scale at which the global Peccei-Quinn (PQ) symmetry $U(1)_{\rm PQ}$ has to be broken in QCD axion models that solve the strong CP problem of the SM [38, 39, 40, 41]. Axion can also be a viable cold dark matter (DM) candidate [42, 43, 44]. Even more generally, axion-like particles (ALPs) are well-motivated, as they naturally appear as pseudo Nambu-Goldstone bosons in many BSM extensions with a spontaneously broken global $U(1)$ symmetry, e.g. in string theory realizations [45, 46, 47], in models of natural inflation [48, 49, 50], baryogenesis [51, 52, 53, 54, 55, 56] and dark energy [57, 58, 59, 60, 61, 62], in the relaxion mechanism for solving the hierarchy problem [63], or in unified ultraviolet-completions of the SM [64, 65, 66, 67, 68]. If the $U(1)$ symmetry breaking happens to be of first-order and strong enough, it can potentially lead to an observable SGWB at current or future GW detectors [69, 70, 71]. In the past years, this aspect raised limited interest because the low-energy ALP phenomenology relevant for laboratory experiments does not depend on the nature of the PT. Today, however, the situation is rather different, since the opportunity of observing GW signals is concrete and offers a uniquely powerful way to directly test the high-scale PQ dynamics, in a way complementary to various laboratory and astrophysical probes of the ALP couplings to the SM particles [72] which go as the inverse of the PQ scale, with model-dependent ${\cal O}(1)$ coefficients [73, 74, 75, 76]. Moreover, the viable parameter space for ALP masses and couplings spans many orders of magnitude, which makes it challenging to completely rule it out; therefore, new ideas and approaches to probe previously inaccessible regions are worth paying attention to. For instance, if the ALP is effectively decoupled from the SM, the only bound may come from black hole superradiance [77]. The GW signal from a FOPT induced by ALPs is another such probe which only depends on the $U(1)$ breaking scale but not on the ALP mass.
@@
During cosmic inflation, primordial overdensities may collapse into PBHs after re-entering the Hubble horizon in the post-inflationary period [126]. However, for single-field inflationary scenarios, the resulting PBH abundance is exponentially sensitive to the amplitude of the curvature perturbations, and therefore it requires extreme fine-tuning [127]. [^2] Fine-tuning may be reduced by introducing poles in the inflaton kinetic term [128] or in the presence of spectator fields during inflation [129]. In this respect, other formation mechanisms such as collapsing of false vacuum bubbles, cosmic strings, domain walls or from preheating [130, 131, 132, 133] may reduce such fine-tuning. In this work, we will focus on PBH formation from supercooled FOPTs. Although this mechanism was first suggested in the pioneering works of Refs. [134, 135], it has recently received great interest due to its successful implementation in BSM scenarios with predictable results that can be tested [136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146].
The core idea is simple: FOPTs proceed via the nucleation of bubbles of the broken phase in an initial background of the symmetric phase [147, 148, 149]. In a supercooled regime, the energy density of the Universe in the symmetric phase is dominated by the vacuum energy which acts as a cosmological constant and leads to an inflationary period. In a nucleated bubble, instead, such energy is quickly converted into radiation and the corresponding patch expands much slower. Since bubble nucleation is a stochastic phenomenon, and since, in a supercooled regime, regions outside the nucleated bubbles expand much faster than those inside, a delayed nucleation within a causal patch can develop a large overdensity that eventually collapses, if large enough, into a PBH. The production mechanism of PBHs by such “late-time blooming” during strong FOPTs, was first studied in Refs. [150, 151, 152, 153, 135, 154], but has recently received a lot of attention [139, 155, 141, 156, 144, 157, 158, 159, 160, 161, 162], also in connection to the PTA signals observed in the form of a SGWB [28, 163, 164, 165]. We will study the formation of PBHs from supercooled FOPT in an explicit ALP realization, where there are only two relevant free parameters, namely, the ALP decay constant $f_{a}$ and a non-Abelian gauge coupling $g$. This makes the model very predictive.
@@
We show that the constraint from the overproduction of PBHs, as well as the bounds on their mass from BBN, CMB and microlensing [166], can strengthen the interplay between GW searches and laboratory, astrophysical and cosmological constraints on the ALP decay constant $f_{a}$ [167]. Through this three-pronged complementarity, we are able to severely cut down the allowed ALP parameter space. Specifically, the GW sensitivity reach of future interferometers and the PBH parameter space ruled out by the aforementioned constraints can be directly mapped onto the ALP parameter space, which provides new bounds on ALP couplings. Conversely, depending on the strength of ALP interactions to SM particles, the bound on the ALP decay constant can be translated into a lower bound on the PBH mass scale. As a side note, we also show that the recent NANOGrav signal of stochastic GW in the nHz frequency range can be explained in our model with $f_{a}\simeq(10~{}{\rm GeV}-1~{}{\rm TeV})$ and $g\simeq 0.56-0.60$; however, in the simplest realization, this falls in the PBH overclosure region ($f_{\rm PBH}>1$).
@@ 2.1 Finite-temperature effects
[…]
@@ 3 GW signals from supercooled ALP phase transition
where $\rho_{{\rm GW}}$ is the energy density in GWs, $\rho_{c}$ is the critical density of the Universe, and $h$ is the dimensionless Hubble parameter. After the GWs are produced in the early Universe they are red-shifted and we observe them today with an amplitude [174]
@@
$h^{2}\Omega_{\rm hyb}(f)=5.10\times 10^{-9}\left(\frac{100}{g_{\ast}(T_{{\rm RH}})}\right)^{1/3}\left(\frac{10}{\beta_{H}}\right)^{2}S_{\rm hyb}(f),$ (3.5)
@@
$\tilde{f}_{\rm hyb}=1.1\times 10^{-4}\;\mathrm{Hz}\,\left(\frac{g_{\ast}}{100}\right)^{1/6}\left(\frac{\beta_{H}}{10}\right)\,\left(\frac{T_{{\rm RH}}}{500\;\mathrm{GeV}}\right).$ (3.7)
The expressions above are only valid for the FOPT we are interested in, that is supercooled PT. This basically means we are considering the limit $\alpha\gg 1$, and ultra-relativistic bubble-wall velocity $v_{w}\simeq 1$.
Other estimates of the GW spectral shapes are available in the literature, e.g. the one obtained using $(3+1)$-dimensional lattice simulations of thick wall bubble collisions [177] and the semi-analytical estimates developed in Refs. [178, 179]. We checked that all these estimates provide similar results after requiring the correct scaling $h^{2}\Omega_{\rm GW}\sim f^{3}$ for super-horizon modes [180, 181, 182, 183, 184].
@@ 4 PBH formatiom from supercooled ALP phase transition
The ratio $\rho_{\rm rad}^{\rm late}/\rho_{\rm rad}^{\rm bkg}$ between the energy densities of the late-blooming patch and of the background keeps increasing until it reaches a maximum value, approximately determined by the percolation temperature of the former. If the overdensity of the late-nucleated patch, quantified by the density contrast $\delta=\rho_{\rm rad}^{\rm late}/\rho_{\rm rad}^{\rm bkg}-1$, is larger than a critical threshold, usually $\delta_{c}\simeq 0.45$, the patch can collapse into a PBH [185] (see Refs. [144, 139, 159, 158] for more details [^7] There are minor differences in estimation of nucleation time between the studies in Ref. [144, 139, 159, 158] based on contribution from past light cones which may lead to slightly varied abundances of PBHs.). We do not go into the details of such computations and instead follow Ref. [159] for our estimates.
@@ 5.1 GW-PBH complementarity
The three purple dashed lines delineate the values of $f_{a}$ and $g$ for which the fraction of DM in PBHs is equal to $f_{\textrm{PBH}}=10^{10},1,10^{-10}$, exemplifying, respectively, an overabundant PBH scenario, one in which the distribution of PBHs can exactly explains the DM relic, and a underabundant PBH case. The $f_{\rm PBH}=1$ case is characterized by values of $\beta/H\sim 5-7$, and the region above it is shown as hatched. In the slice of the plot defined by $110^{-10}<f_{\textrm{PBH}}<$ 1, we report the constraints, given by shaded cyan regions, from the PBH searches that we will be detailed below. Concurrently, we highlight in red, on top of the $f_{\textrm{PBH}}=1$ curve, the PBH mass window compatible with all the PBH bounds (see Section 5.4). This is realized by $\beta/H\sim 6.5-7$. See Fig. 4 for a zoomed-in view of this region.
@@
Figure 3: PBH and GW predictions in our scale-invariant $SU(2)$ realization of the ALP model. Some contours of $f_{\rm PBH}$ (purple), $\alpha$ and $\beta/H$ are shown. The gray and red shaded regions are disfavored. The thick red line shows the allowed region where PBHs can explain 100% DM relic. The blue shaded region is the $2\sigma$ preferred region to explain the NANOGrav signal. Other solid/dashed lines correspond to the sensitivities of the future GW detectors. See text for further details.
@@
where $\Omega_{\rm th}h^{2}$ and $\Omega_{\rm exp}h^{2}$ represent, respectively, the GW relic from our theoretical prediction of the supercooled FOPT and the experimental value from NANOGrav. For the estimate of the $\chi^{2}$ we ignore the uncertainty in the frequency width and just consider, for each bin, the uncertainty $\sigma_{i}$ in the $\log_{10}(\Omega_{\rm exp}h^{2})$, which we approximate using the half-width. We see that the interpretation of the NANOGrav signal as a stochastic background of GW from a supercooled FOPT in our ALP model is disfavored since it falls within the PBH overclosure region. However, this is only true if we assume instantaneous reheating. If there is a preheating or prolonged reheating period, aided by PBH evaporation (see e.g. Ref. [209]), some (or most) of the PBH population can in principle be diluted away by the end of reheating to relax the overclosure bound and save the NANOGrav-preferred region. We postpone this discussion to a future work.
@@
[…]
@@
[…]

### v2

@@ Abstract
We study the formation of primordial black holes (PBHs) from density fluctuations due to supercooled phase transitions (PTs) triggered in an axion-like particle (ALP) model. We find that the mass of the PBHs is inversely correlated with the ALP decay constant $f_{a}$. For instance, for $f_{a}$ varying from $\cal O$}(100 MeV) to $\cal O$}($10^{12}$ GeV), the PBH mass varies between $(10^{3}-10^{-24})M_{\odot}$. We then identify the ALP parameter space where the PBH can account for the entire (or partial) dark matter fraction of the Universe, in a single (multi-component) dark matter scenario, with the ALP being the other dark matter candidate. The PBH parameter space ruled out by current cosmological and microlensing observations can thus be directly mapped onto the ALP parameter space, thus providing new bounds on ALPs, complementary to the laboratory and astrophysical ALP constraints. Similarly, depending on the ALP couplings to other Standard Model particles, the ALP constraints on $f_{a}$ can be translated into a lower bound on the PBH mass scale. Moreover, the supercooled PT leads to a potentially observable stochastic gravitational wave (GW) signal at future GW observatories, such as aLIGO, LISA and ET, that acts as another complementary probe of the ALPs, as well as of the PBH dark matter. Finally, we show that the recent NANOGrav signal of stochastic GW in the nHz frequency range can be explained in our model with $f_{a}\simeq(0.1~{\rm GeV}-1~{\rm TeV})$.
@@ 1 Introduction
[…]
@@
Interestingly, the FOPT energy scale LIGO is sensitive to roughly coincides with the lowest possible energy scale at which the global Peccei-Quinn (PQ) symmetry $U(1)_{\rm PQ}$ has to be broken in QCD axion models that solve the strong CP problem of the SM [38, 39, 40, 41]. Axion can also be a viable cold dark matter (DM) candidate [42, 43, 44]. Even more generally, axion-like particles (ALPs) are well-motivated, as they naturally appear as pseudo Nambu-Goldstone bosons in many BSM extensions with a spontaneously broken global $U(1)$ symmetry, e.g. in string theory realizations [45, 46, 47], in models of natural inflation [48, 49, 50], baryogenesis [51, 52, 53, 54, 55, 56] and dark energy [57, 58, 59, 60, 61, 62], in the relaxion mechanism for solving the hierarchy problem [63], or in unified ultraviolet-completions of the SM [64, 65, 66, 67, 68]. If the $U(1)$ symmetry breaking happens to be of first-order and strong enough, it can potentially lead to an observable SGWB at current or future GW detectors [69, 70, 71, 72]. In the past years, this aspect raised limited interest because the low-energy ALP phenomenology relevant for laboratory experiments does not depend on the nature of the PT. Today, however, the situation is rather different, since the opportunity of observing GW signals is concrete and offers a uniquely powerful way to directly test the high-scale PQ dynamics, in a way complementary to various laboratory and astrophysical probes of the ALP couplings to the SM particles [73] which go as the inverse of the PQ scale, with model-dependent ${\cal O}(1)$ coefficients [74, 75, 76, 77]. Moreover, the viable parameter space for ALP masses and couplings spans many orders of magnitude, which makes it challenging to completely rule it out; therefore, new ideas and approaches to probe previously inaccessible regions are worth paying attention to. For instance, if the ALP is effectively decoupled from the SM, the only bound may come from black hole superradiance [78]. The GW signal from a FOPT induced by ALPs is another such probe which only depends on the $U(1)$ breaking scale but not on the ALP mass.
@@
During cosmic inflation, primordial overdensities may collapse into PBHs after re-entering the Hubble horizon in the post-inflationary period [126]. However, for single-field inflationary scenarios, the resulting PBH abundance is exponentially sensitive to the amplitude of the curvature perturbations, and therefore it requires extreme fine-tuning [127]. [^2] Fine-tuning may be reduced by introducing poles in the inflaton kinetic term [128] or in the presence of spectator fields during inflation [129]. In this respect, other formation mechanisms such as collapsing of false vacuum bubbles, cosmic strings, domain walls or from preheating [130, 131, 132, 133, 134, 135] may reduce such fine-tuning. In this work, we will focus on PBH formation from supercooled FOPTs. Although this mechanism was first suggested in the pioneering works of Refs. [136, 137], it has recently received great interest due to its successful implementation in BSM scenarios with predictable results that can be tested [138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148].
The core idea is simple: FOPTs proceed via the nucleation of bubbles of the broken phase in an initial background of the symmetric phase [149, 150, 151]. In a supercooled regime, the energy density of the Universe in the symmetric phase is dominated by the vacuum energy which acts as a cosmological constant and leads to an inflationary period. In a nucleated bubble, instead, such energy is quickly converted into radiation and the corresponding patch expands much slower. Since bubble nucleation is a stochastic phenomenon, and since, in a supercooled regime, regions outside the nucleated bubbles expand much faster than those inside, a delayed nucleation within a causal patch can develop a large overdensity that eventually collapses, if large enough, into a PBH. The production mechanism of PBHs by such “late-time blooming” during strong FOPTs, was first studied in Refs. [152, 153, 154, 155, 137, 156], but has recently received a lot of attention [141, 157, 143, 158, 146, 159, 160, 161, 162, 163, 164, 165], also in connection to the PTA signals observed in the form of a SGWB [28, 166, 167, 168]. We will study the formation of PBHs from supercooled FOPT in an explicit ALP realization, where there are only two relevant free parameters, namely, the ALP decay constant $f_{a}$ and a non-Abelian gauge coupling $g$. This makes the model very predictive.
@@
We show that the constraint from the overproduction of PBHs, as well as the bounds on their mass from BBN, CMB and microlensing [169], can strengthen the interplay between GW searches and laboratory, astrophysical and cosmological constraints on the ALP decay constant $f_{a}$ [170]. Through this three-pronged complementarity, we are able to severely cut down the allowed ALP parameter space. Specifically, the GW sensitivity reach of future interferometers and the PBH parameter space ruled out by the aforementioned constraints can be directly mapped onto the ALP parameter space, which provides new bounds on ALP couplings. Conversely, depending on the strength of ALP interactions to SM particles, the bound on the ALP decay constant can be translated into a lower bound on the PBH mass scale. As a side note, we also show that the recent NANOGrav signal of stochastic GW in the nHz frequency range can be explained in our model with $f_{a}\simeq(0.1~{\rm GeV}-1~{\rm TeV})$ while satisfying the PBH constraints.
@@ 2.1 Finite-temperature effects
$J_{B/F}(y^{2})=\int_{0}^{\infty}dt~t^{2}\log{\left[1\mp\exp{\left(-\sqrt{t^{2}+y^{2}}\right)}\right]}\,.$ (2.4)
@@ 3 GW signals from supercooled ALP phase transition
where $\rho_{{\rm GW}}$ is the energy density in GWs, $\rho_{c}$ is the critical density of the Universe, and $h$ is the dimensionless Hubble parameter [^6] The generation of Gravitational Waves occur during the bubble collision, sourced by the non-zero anisotropic stress component in their space-space part $\Pi_{ij}$, see Refs. [31, 177, 178] for a review.. After the GWs are produced in the early Universe they are red-shifted and we observe them today with an amplitude [179]
@@
$h^{2}\Omega_{\rm hyb}(f)=5.10\times 10^{-9}\left(\frac{100}{g_{\ast}(T_{{\rm RH}})}\right)^{1/3}\left(\frac{10}{\beta/H}\right)^{2}S_{\rm hyb}(f),$ (3.5)
@@
$\tilde{f}_{\rm hyb}=1.1\times 10^{-4}\;\mathrm{Hz}\,\left(\frac{g_{\ast}}{100}\right)^{1/6}\left(\frac{\beta/H}{10}\right)\,\left(\frac{T_{{\rm RH}}}{500\;\mathrm{GeV}}\right).$ (3.7)
The expressions above are only valid for the FOPT we are interested in, that is supercooled PT. This basically means we are considering the limit $\alpha\gg 1$, and ultra-relativistic bubble-wall velocity $v_{w}\simeq 1$. We also enforce the correct scaling $h^{2}\Omega_{\rm GW}\sim f^{3}$ for super-horizon modes which correspond to the frequencies at IR tail of the spectrum below [182, 183, 184, 185, 186]
$f_{\rm horizon}=0.013\,{\rm mHz}\left(\frac{g_{\ast}}{100}\right)^{1/6}\left(\frac{T_{{\rm RH}}}{500\;\mathrm{GeV}}\right)\,.$ (3.8)
Other estimates of the GW spectral shapes are available in the literature, e.g. the one obtained using $(3+1)$-dimensional lattice simulations of thick wall bubble collisions [187] and the semi-analytical estimates developed in Refs. [188, 189]. We checked that all these estimates provide similar results after requiring the correct scaling $h^{2}\Omega_{\rm GW}\sim f^{3}$ for super-horizon modes.
@@ 4 PBH formatiom from supercooled ALP phase transition
The ratio $\rho_{\rm rad}^{\rm late}/\rho_{\rm rad}^{\rm bkg}$ between the energy densities of the late-blooming patch and of the background keeps increasing until it reaches a maximum value, approximately determined by the percolation temperature of the former. If the overdensity of the late-nucleated patch, quantified by the density contrast $\delta=\rho_{\rm rad}^{\rm late}/\rho_{\rm rad}^{\rm bkg}-1$, is larger than a critical threshold, usually $\delta_{c}\simeq 0.45$, the patch can collapse into a PBH [190] (see Refs. [146, 141, 161, 160] for more details [^8] There are minor differences in estimation of nucleation time between the studies in Ref. [146, 141, 161, 160] based on contribution from past light cones which may lead to slightly varied abundances of PBHs.). We do not go into the details of such computations and instead follow Ref. [161] for our estimates [^9] We follow the analysis of this section, satisfying Hoop’s conjecture for PBH formation as discussed in Refs. [160, 191]..
@@ 5.1 GW-PBH complementarity
The three purple dashed lines delineate the values of $f_{a}$ and $g$ for which the fraction of DM in PBHs is equal to $f_{\textrm{PBH}}=10^{10},1,10^{-10}$, exemplifying, respectively, an overabundant PBH scenario, one in which the distribution of PBHs can exactly explains the DM relic, and a underabundant PBH case. The $f_{\rm PBH}=1$ case is characterized by values of $\beta/H\sim 5-7$, and the region above it is shown as hatched. In the slice of the plot defined by $10^{-10}<f_{\textrm{PBH}}<1$, we report the constraints, given by shaded cyan regions, from the PBH searches that we will be detailed below. Concurrently, we highlight in red, on top of the $f_{\textrm{PBH}}=1$ curve, the PBH mass window compatible with all the PBH bounds (see Section 5.4). This is realized by $\beta/H\sim 6.5-7$. See Fig. 4 for a zoomed-in view of this region.
@@
Figure 3: PBH and GW predictions in our scale-invariant $SU(2)$ realization of the ALP model. Some contours of $f_{\rm PBH}$ (purple), $\alpha$ and $\beta/H$ are shown. The gray and red shaded regions are disfavored. The thick red line shows the allowed region where PBHs can explain 100% DM relic. The blue shaded regions are the $2\sigma$ preferred regions to explain the NANOGrav signal. Other solid/dashed lines correspond to the sensitivities of the future GW detectors. See text for further details.
@@
[…]
@@
$\mathrm{SNR}=\sqrt{t_{\rm obs}\int_{f_{\rm min}}^{f_{\rm max}}df\bigg[\frac{\Omega(f)}{\Omega_{\rm noise}(f)}\bigg]^{2}}\,,\qquad\Omega_{\rm noise}(f_{\rm gw})=\frac{2\pi^{2}}{3H_{0}^{2}}f^{3}S_{\rm noise}(f_{\rm gw})\,$, (5.4)
@@
$=$ $\frac{\mathrm{SNR}_{\rm ref}}{\sqrt{t_{\rm obs}}}\max_{b}\bigg[\bigg(\int_{f_{\rm min}}^{f_{\rm max}}d\bar{f}\frac{\bar{f}^{2b}}{\Omega_{\rm noise}^{2}(\bar{f})}\bigg)^{-\frac{1}{2}}\,f_{\rm gw}^{b}\bigg]\,.$